OpenStudy (anonymous):

Function table Help please?? Urgent. I need to complete it, and I have know one to help. Function rule is?? Input Output -3 ? -2 -4 -1 ? 0 ? 1 -1 _________________________________ I tried tried and tried and brain is blocked. Just send me in the right direction? I can't stay up much longer. have to go to bed soon.

OpenStudy (anonymous):

It looks like the output is just equal to the input minus 2. Do you see it?

OpenStudy (anonymous):

no so -2 -3??

OpenStudy (anonymous):

Take the input, and subtract 2 from it. That will give you the output. For example, take the -2 input. (-2)-2 = -4

OpenStudy (anonymous):

(-3) -2 = -5 ?

OpenStudy (anonymous):

Mhmm, you got it!

OpenStudy (primeralph):

........could be quadratic...or exponential..........or logarithmic...........

OpenStudy (anonymous):

oh thank you so much! Did I write the function rule right? y=-2

OpenStudy (anonymous):

You probably need to solve the integral using Green's Theorem too

OpenStudy (primeralph):

integral?

OpenStudy (anonymous):

And no, x is the input, while y is the output. y=x-2

OpenStudy (anonymous):

I was just joking along with you Ralph ^^

OpenStudy (anonymous):

ok thank you @smokeydabear your awesome!

OpenStudy (anonymous):

Yeah no problem, glad I could help!